You are considering a preferred stock for purchase. The stock has a par value of $22 and, based on the risk of the stock you want to earn a required return of 5.2%. If the stock pays a 10.9% fixed dividend, your estimate of the price of the stock would be $__.__. Round your answer to two decimal places.
Answer #1

Preferred dividend = 0.109 * 22 = 2.398

Price of stock = Perpetual dividend / required return

Price of stock = 2.398 / 0.052

Price of stock = $46.12